OpenStudy (abbyrose0105):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3? Simplify the given expression to rational exponent form and justify each step by identifying the properties of rational exponents used. All work must be shown. 1 over the cube root of the quantity of x to the negative 6 pwr Simplify the given expression to radical form and justify each step by identifying the properties of rational exponents

OpenStudy (neer2890):

x3 mean 3x or\[x ^{3}???\]

OpenStudy (abbyrose0105):

Thanks!!

OpenStudy (neer2890):

for what...?

OpenStudy (abbyrose0105):

I know you didn't actually figure it out but when you wrote that it reminded me of something and I've got it now

OpenStudy (neer2890):

okay. so what's your answer now?

OpenStudy (abbyrose0105):

x^3 times x^3 times x^3 times x^3 = x^12 and x^3+x^3+x^3+x^3=4x^3

OpenStudy (abbyrose0105):

I still don't have an answer to the second one though Simplify the given expression to rational exponent form and justify each step by identifying the properties of rational exponents used. All work must be shown. 1 over the cube root of the quantity of x to the negative 6 pwr

OpenStudy (neer2890):

you mean you have to solve this\[\frac{ 1 }{ \sqrt[3]{x ^{-6}} }????\]

OpenStudy (abbyrose0105):

yes

OpenStudy (neer2890):

\[=\frac{ 1 }{ \sqrt[3]{\frac{ 1 }{ x ^{6} }} }=\frac{ 1 }{ \sqrt[3]{(\frac{ 1 }{ x ^{2} })^{3}} }=\frac{ 1 }{ \frac{ 1 }{ x ^{2} } }=x ^{2}\]

OpenStudy (abbyrose0105):

How do I solve that? I want to understand it so I can do other problems similar to this

OpenStudy (neer2890):

ok. do you know that \[x ^{-6}=\frac{ 1 }{ x ^{6} }??\]

OpenStudy (abbyrose0105):

yes

OpenStudy (neer2890):

and we can write \[x ^{6}=(x ^{2})^{3}\] because we have to eliminate the cube root, that is why we have to break 6 into such power that it contains 3. and power 3 and cube root cancels with each other.

OpenStudy (neer2890):

did you get it...? take your time

OpenStudy (abbyrose0105):

yes I think that makes sense

OpenStudy (neer2890):

so,you need any other help or you can do it now....?

OpenStudy (abbyrose0105):

is the answer x^2?

OpenStudy (neer2890):

correct...! good work..!